OpenStudy (anonymous):

solve -3 - { -9 - [ -7 - (6 - 12 ]}

OpenStudy (unklerhaukus):

start with simplfying the inner most brackets

OpenStudy (unklerhaukus):

(6-12)

OpenStudy (unklerhaukus):

can you simplify six minus twelve?

OpenStudy (anonymous):

-6

OpenStudy (unklerhaukus):

cool, so now you have; -3 - { -9 - [ -7 - (-6)]} can you simplify this bit [ -7 - (-6)] ?

OpenStudy (anonymous):

1. 6-12 = -6 2. -7 - (-6) = -1 3. -9 - (-1) = -8 4. -3 - (-8) = 5

OpenStudy (anonymous):

I think 5 is the answer

OpenStudy (unklerhaukus):

All those minus signs make it confusing , but you haven't made a mistake, well done, 5 is correct \(\checkmark\) @marcyabc123

OpenStudy (anonymous):

Thank you!